OpenStudy (anonymous):

x^2+4x-4y+16=0 -put in standard form -find vertex -find focus

OpenStudy (campbell_st):

then 4y = x^2 +4x + 16 complete the square for x 4y = (x + 2)^2 + 12 4y - 12 = (x+2)^2 4(y - 3) = (x +2)^2 now in the standard form vertex is at (-2, 3)

OpenStudy (campbell_st):

x^2 = 4ay in this question (x+2)^2 = 4(y-3) so a= 1 focus is at (-2, 4)

OpenStudy (anonymous):

so whats the final answer in standard form?

OpenStudy (campbell_st):

which standard form \[(x - h)^2 = 4a(y-k) \] or \[y =ax^2+bx + c\]

OpenStudy (campbell_st):

and the vertex and focus don't have a standard form